Recently, the Transport Association of Zambia increased the price of bus fares for all long routes in Zambia. Specifically, the price of tickets between Lusaka and Livingstone increased from 200 Kwacha to

250 kwacha. The demand equation between Lusaka and Livingstone.

a) Calculate the old price quantity demanded level

Marks]

b) Calculate the new price quantity demanded level

c) Work out the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200

kwacha and 250 kwacha [5

Marks]

d) Briefly discuss the Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha and clarify how bus fleets revenues are expected to respond to the price increase. [5 Marks]

e) What would be the Price Elasticity of Demand for tickets from Lusaka to Livingstone if the price rises again from 250 kwacha to 300 kwacha? [5 Marks]

f) Explain why the answer in part d) and part e) are different.


1
Expert's answer
2021-09-09T17:20:39-0400

Qd=2000-5P

A. When P=200 kwacha, "Qd=2000-5 \times 200=1000"

Old price quantity demanded=1000

B. When P=250 kwacha, "Qd=2000-5 \times 250=750"

New price quantity demanded=1500.

C. Price elasticity of demand

"=\frac{(\frac{(Q2-Q1)}{(Q1+Q2)/2})}{(\frac{(P2-P1)}{(P2+P1)/2})}"

"= \frac{(\frac{(750-1000)}{(750+1000)/2})}{(\frac{(250-200)}{(250+200)/2}) } \\

=-1.3"

D.

Price Elasticity of Demand for tickets from Lusaka to Livingstone between 200 kwacha and 250 kwacha is elastic. Since the absolute value of elasticity is more than 1. (1.3>1).

So rise in price will lead to fall in total revenue. As rise in price will lead to more fall in quantity demanded. So total revenue will fall.

E.

When P=300 kwacha, "Qd=2000-5 \times 300=500"

Price elasticity of demand "=\frac{(\frac{(Q2-Q1)}{(Q1+Q2)/2})}{(\frac{(P2-P1)}{(P2+P1)/2})}"

"= \frac{(\frac{(750-500)}{(750+500)/2})}{(\frac{(250-300)}{(250+300)/2}) } \\

=-2.2"

F. When price rises from 250 to 300 kwacha, demand is still elastic buy it is more elastic than when price rises from 200 to 250.

So total revenue will still fall fue to rise in prices.